OpenStudy (anonymous):

Find an equation of the line passing through the given points. Use function notation to write the equation. (1,-3) and (2,-1) f(x)=

OpenStudy (micahm):

ok so this is what i going to do (1,-3) and (2,-1) f(x)= is

OpenStudy (micahm):

i this how it is written

OpenStudy (anonymous):

Is that the answer? @micahm

OpenStudy (micahm):

no is that how it is written

OpenStudy (micahm):

does this have a abcd chose

OpenStudy (micahm):

If we multiply both sides of this last equation by x − x0, we get ... Chapter 3 Linear Functions ... 2. Substitute the coordinates of the given point for x0 and y0 in the formula ... For example, if the line has slope −2 and passes through the point (3,4), then ... m = 1/2. Use the point-slope form to determine the equation of the line.

OpenStudy (anonymous):

So (1/3) (2x--1) Right?

OpenStudy (micahm):

yeah

OpenStudy (anonymous):

Then how do you finish it? @micahm

OpenStudy (micahm):

nd the equation of the line passing through the given points. write the equation in function notation. (2,-8) and (-4,-3) ** Equation of a line: ... =(-8-(-3))/(2-(-4)) =(-8+3)/(2+4) m=-5/6. Equation: f(x)=-5x/6+b solve for b using coordinates of one of given points (2,-8) -8=-5*2/6+b -48=-10+6b 6b=-38 b=-38/6=-19/3. Equation:

OpenStudy (anonymous):

Woah you lost me there @micahm

OpenStudy (micahm):

sorry about that

OpenStudy (anonymous):

Can you explain it again @micahm

OpenStudy (micahm):

(1,-3) and (2,-1) f(x)= Find an equation do you know what an equation is first of all

OpenStudy (anonymous):

Yes I do

OpenStudy (micahm):

which is

OpenStudy (anonymous):

a formula

OpenStudy (micahm):

this is not graphing is it

OpenStudy (micahm):

are you there

OpenStudy (micahm):

hello

OpenStudy (anonymous):

hi sorry

OpenStudy (anonymous):

not its not graphing

OpenStudy (micahm):

ok have you ever hared of http://www.freemathhelp.com/

OpenStudy (micahm):

have you heard of it

OpenStudy (micahm):

are you there

OpenStudy (anonymous):

yes

OpenStudy (mrnood):

This is exactly the same as the last questions you had help with. You have not made any effort to solve those for yourself. First do exactly as I tried to show you before find (y2-y1) find (x2-x1) find slope = (y2-y1)/(x2-x1) If you just wait for someone else to give you the answer you will never understand how to do these questions

OpenStudy (anonymous):

I suck at math though

OpenStudy (mrnood):

Would you proudly say 'I suck at English' or 'I suck at .....'?? Maths is just a skill to be learned - it is nor more difficult than geography, French or other subject if you address the learning process an not be scared by the process. You have asked at least 3 identical questions today and not made a single positive contribution towards solving them. A straight line is a fairly simple thing and you should not be scared off by the suffixes and equations. The word 'slope' means the same as it does in English - how rapidly does a line (or hill or ramp or....) change. If you have 2 known points then you can draw a line between them - its slope is defined by the points and is the same for all points on the line. If you know 1 point and the slope then you can draw the line through the point. In many parts of maths you will find that IF you can draw or graph the problem then the Math is just a way of doing the same with out a pencil. See if you can draw the line in this question and mark on it the 2 points and the slope.... it will help you overcome the 'unknown' where to start and might help you believe that you only 'suck at Math' because you think you do.